solve by using the quadratic formula?
x^2=-5x+8

Answer by checkley75(3666)

x^2=-5x+8
x^2+5x-8=0
x=(-5+-sqrt[5^2-4*1*-8)]/2*1
x=(-5+-sqrt[25+32])/2
x=(-5+-sqrt57)/2
x=(-5+-7.55)/2
x=(-5+7.55)/2
x=2.55/2
x=1.275 answer.
x=(-5-7.55)/2
x=-12.55/2
x=-6.275 answer.